# Space science & the space economy

### v1

@@ Abstract
Will it be possible in the future to realize large, complex space missions dedicated to basic science like HST, Chandra and JWST? Or will their cost be just too great? Today’s space scene is completely different from that of even five years ago, and certainly from that of the time when HST, Chandra and JWST were conceived and built. Space-related investments have grown exponentially in recent years, with monetary investment exceeding half a trillion dollars in 2023. This boom is largely due to the rise of the so-called ’new space’ economy driven by private commercial funding, which for the first time last year surpassed public investments in space. The introduction of a market logic to space activities results in more competition, and a resulting dramatic cost and schedule reduction. Can space science take advantage of the benefits of the new space economy to reduce cost and development time and at the same time succeed in producing powerful missions in basic science? The prospects for Europe and the USA are considered here. We argue that this goal would be attainable if the scientific community could take advantage of the three pillars underlying the innovation of the new space economy: (1) technology innovation proceeding through both incremental innovation and disruptive innovation, (2) business innovation, through vertical integration, scale production, and service oriented business model, and (3) cultural innovation, through openness to risk and iterative development.
@@ 1 Basic science from space
    multimessenger astrophysics
@@
The first was still beyond the horizon when the James Webb Space Telescope (JWST, at the time called Next Generation Space Telescope, NGST) was recommended at the end of the 1990s as the highest priority in the 2000 Decadal Survey. The other three were in their infancy, with the first discoveries just announced. They were not among the main science drivers for NGST. Instead, in the Decadal report ’Astronomy and Astrophysics in the New Millennium’ [1] one reads that NGST ’..is designed to detect light from the first stars and to trace the evolution of galaxies from their formation to the present. It will revolutionize the understanding of how stars and planets form in our galaxy today.. thanks to Kuiper Belt objects (KBOs) in our solar system, the formation of stars and planets in our galaxy, and the dust emission from galaxies out to redshifts of 3’. JWST is doing superb work on the earliest galaxies. Yet JWST is doing a great job on all four of the new themes: up to 40% of JWST time is dedicated to exoplanet studies alone. JWST can be this flexible because it was designed to be a multipurpose observatory that pushes the sensitivity for both photometry and spectroscopy by orders of magnitude with respect to previous or contemporary facilities, opening up new ’discovery space.’ This powerful capability did not come for free. The complexity, cost and development time of JWST are large, even huge, with more than 300 single-point failure modes [2], a cost of about 9B\$ and 23 years elapsed from the publication of the 2000 decadal survey to launch. This cost should be compared with the predicted cost to completion for NGST in ’Astronomy and Astrophysics in the New Millennium’, given as 1B\$. This almost 10-fold cost growth of NGST cannibalized other NASA Astrophysics programs, notably the ’Beyond Einstein’ program spanning a wide range of other wavelengths [3]. Concentration on NGST introduced an enormous risk to space-based astrophysics by putting all its eggs in one basket: a failure of JWST would have meant a decades-long setback for all observational astrophysics.
@@
Even more worrisome is that it is far from certain that NASA, ESA, or any other space agency will be willing or able to embark on a new JWST-scale project (i.e., \$10B, $>20$ year development time). In April 2024 NASA paused the original JPL-led Mars Sample Return project [^2] https://www.nasa.gov/news-release/nasa-sets-path-to-return-mars-samples-seeks-innovative-designs/, which was in the ballpark cost and development time range as JWST. NASA issued a call to industry for alternative approaches and in June 2024 selected seven companies to conduct studies of ’out of the box’ concepts that can deliver samples faster and cheaper than the current plan [^3] https://www.nasa.gov/news-release/nasa-exploring-alternative-mars-sample-return-methods/. This re-thought mission may become another example in which NASA buys a service for science-related or exploration-related missions, as NASA now does in other areas, for example transportation to and from the International Space Station and the Commercial Lunar Payload Service. Another example of the NASA change in attitude is the end of the Moon VIPER rover project [^4] https://www.nasa.gov/news-release/nasa-ends-viper-project-continues-moon-exploration/ due to increases in costs and delays. An additional cause for concern to astrophysics is that NASA’s major effort today and for the next few years will certainly be the Artemis project, returning humans to the lunar surface in a long-term program. How can space astrophysics thrive in these circumstances? Can projects like HWO and Artemis coexist? An even greater element of uncertainty is to predict at this time what the attitude of NASA and the new U.S. government in general toward basic science missions might be.
@@
An alternative to expensive multipurpose observatories is to have multiple dedicated missions. ESA’s Gaia mission, for example, opened up new discovery space by measuring the positions and motions of a billion stars some 1000X better than was possible before. Gaia is becoming the most productive telescope in terms of the number of papers produced from its data, about 2000 papers published per year in refereed journals [^5] https://www.cosmos.esa.int/web/gaia/peer-reviewed-journals, compared with the roughly 1000 papers per year based on HST data [^6] https://science.nasa.gov/mission/hubble/overview/hubble-by-the-numbers/. The ESA Euclid and NASA Nancy Grace Roman telescopes are other examples. The total cost of Gaia and Euclid (ESA plus member state contributions) is in the 1-1.5B\$ each, while that of the Roman telescope is close to 5 B\$ (adding a coronagraph to Roman increased its cost by \$ 350-500M [^7] https://roman.gsfc.nasa.gov/science/rsig/2021/Roman_Requirements_20201105.pdf). Their development time is in the 10-15 years range. That is some 2-10X cheaper and 2X faster than JWST and HWO. Although this is an improvement, even at these costs, it is not clear that space agencies can afford more than one per decade of these missions. Such a pace creates the risk that many emerging scientific topics, and so sectors of the astrophysical community, are left unserved at the margin.
@@
Unlike ESA, NASA has a strong Explorer program [^8] https://explorers.gsfc.nasa.gov/ which provides affordable and frequent flight opportunities for smaller science missions [5]. This ensures ’rapid response to changing science and technology to enable cutting-edge science at moderate cost’. During the long JWST development time, the Explorer program has contributed to maintaining a reasonable level of diversity in the astrophysical ecosystem, avoiding the drying up of entire scientific communities. In Europe, the lack of an Explorer-like program has challenged the high-energy, mid/far-infrared and millimeter communities, which were unable to benefit from replacements for the highly successful XMM-Newton, Herschel and Planck X-ray, far-infrared and CMB- millimeter missions. Of course, Explorer missions are somewhat limited in scope, being dedicated to well-focused scientific goals and in some cases representing case studies and in-orbit demonstrations for new techniques. As an example, the latest SMEX astrophysics mission is IXPE, dedicated to X-ray polarimetry, certainly a niche science. Although IXPE is certainly a highly successful mission (the team was awarded the 2024 Bruno Rossi Prize), its limited collecting area allows sensitive observations of only a few dozen bright sources. IXPE has shown the high potential of X-ray polarimetry, but to fully harvest its fruit a much larger X-ray polarimetric telescope is needed. The current cost cap of Explorer missions is around \$120M - \$130M for the SMEX missions and about \$300M for MIDEX missions999\$300M in NASA Fiscal Year (FY) 2022 dollars: https://nspires.nasaprs.com/external/viewrepositorydocument/cmdocumentid=841540/solicitationId=%7B0F29B64B-AFF1-ADAD-7302-75D7AEF2C696%7D/viewSolicitationDocument=1/MIDEX_AO_Amend%201.pdf. From 2000 seven SMEX and five MIDEX astrophysics missions were launched [^10] https://en.wikipedia.org/wiki/Explorers_Program#Small_Explorers_(SMEX), that is about a mission every two years. The typical development time of Explorer missions is 4-5 years, and the entry-level TRL [^11] TRL = technical readiness level: https://en.wikipedia.org/wiki/Technology_readiness_level is 5-6.
@@ 2 The new space economy
At the same time that JWST was launched and HWO was being conceived, we witnessed the booming of the so-called Space Economy. The total amount of money invested in the Space Economy in 2023 was \$570B according to The Space Report of the Space Foundation [^12] https://www.spacefoundation.org/2024/07/18/the-space-report-2024-q2/, with a growth of 7.4% with respect to 2022, for a doubling time of less than 10 years. Recent years have seen private funding overtake public funding. Growth may well accelerate in the next few years, reaching the threshold of \$1000B at the end of this decade and up to 1.8 trillion dollars by 2035131313https://www3.weforum.org/docs/WEF_Space_2024.pdf, dwarfing any conceivable increase in government space agency budgets (except maybe military). Many of these start-ups go well beyond providing launch into on-orbit applications such as Internet communications and many forms of Earth observation. Satellite numbers have grown from less than 200/year in 2000, to almost 3000/year in 2023141414https://www.planet4589.org/space/stats/pay.html .
@@
Can the scientific community take advantage of the new space economy to produce ambitious new space missions dedicated to basic science at a much lower cost and faster timeline? To investigate this issue, we need to better understand the differences between the old space economy and the new one. Table 1 provides a useful simplification as a first approach.
@@
Future: The public-funded space economy will more than double by 2035. Expenditure on defense may triple. Other areas of great increase: PNT; infrastructures. New areas of investment: manufacturing in space, including pharmaceutics, biotechnology and semiconductors; space mining; space solar production; space travel; In the past decades activity/business started in the old space-economy framework using public money and then progressed to the new space economy. By 2035 the new space economy may more than triple. Areas of large increases will probably be: PNT, which alone may consist of $>$75% of all business, for example last mile deliveries and point to point transportation; communications, Earth observation
@@
Another important space resource will be materials [8]: water, oxygen, hydrogen, construction materials including iron, titanium, and rock, and precious metals, including the Platinum group elements (PGE). Rare Earth elements (REE) are also common on the Moon in the KREEP terrane [9]. REE were subject to large price volatility in the past decades, in particular when China put strict export quotas in 2009-2010 [10, 11]. Most of the materials extracted would be directly used in space. Only in the case of precious metals or strategic elements would they be brought back on Earth.
The final important resource commonly used in space, but not often acknowledged, is ”knowledge”. Knowledge is crucial to enable or make more efficient harvesting of other resources and to protect space systems. Navigation in low Earth orbits requires detailed knowledge of the Earth’s gravitational field; likewise for the Moon. Navigation near small bodies requires accurate knowledge of their gravitational field and response to solar illumination. Without a deep understanding of the physics of the solar wind and without real-time data it is impossible to forecast solar storms that can easily damage space infrastructures. And, it is impossible to mine for resources the Moon or Mars or an asteroid if it is not known with precision what is present on the surface or interiors of these bodies through prospecting. These are just a few examples, but it is clear that without long and accurate observations, big data analysis, and the development of accurate models it is impossible to imagine any activity in space. Knowledge therefore has its own very concrete value as intellectual property (IP) in the new space economy. This is a major change from the open data policies of most government space agencies.
@@ 2.1 Innovation paradigms
    Innovation in technology, proceeding through both incremental innovation and disruptive innovation. Disruptive innovation starts with low performance and high cost, but quickly overtakes mainstream technologies. Disruptors then engage in incremental innovation, beginning a new cycle.
@@
    Cultural innovation. Startups must have a bold-risk culture, which manifests in risk openness project management models, for example the Agile methodology, which breaks projects into several “sprints” and delivers a Minimum Viable Product (MVP) at the end of each sprint.
@@ 2.2 Costs: launch
Figure 1, adapted from Visual Capitalist, shows the cost per kg to orbit as a function of time. Thanks to reusability, the SpaceX Falcon 9 (F9) and Falcon Heavy (FH) launchers have lowered the cost to orbit offered to third parties by a factor of 3-6 with respect to other Western, Russian, and Chinese launchers. F9 capacity to low Earth orbit (LEO) is about 20 tons, at an advertised cost of about 62M\$, that is about \$3000/kg.
@@
F9 has today an effective monopoly of the launch market. Forthcoming competition with Blue Origin New Glenn may result in a lowered cost for third parties. Table 3 provides examples of launch costs for recent scientific missions. None approach the \$1-3k/kg that is achievable with F9, mainly because they were unable to make use of the full F9 capacity on dedicated launches, especially for missions not designed for an F9 or FH launch. Other causes of high-billed costs include requests for orbits more demanding than the LEO and absence of competition.
@@
Mission Weight kg Orbit Launcher Cost M\$ Cost k\$/kg
@@
The launch cost of Figure 1 has been reached for SpaceX Starlink satellites that have been estimated at 0.7-1.4 M\$ per Starlink 2 mini satellite, with a cost per kg as low as \$1000-\$2000/kg. This should be regarded as the cost of launch for optimized payloads, net of profit. SpaceX claims that the coming on-line of Starship in the next few years will lower the cost by another factor of 10 with respect to F9. The Starship revolution will not be limited to cost to orbit. Starship performance will make it unnecessary to spend time and resources to minimize mass and size [6, 12, 13, 14], which will change the way not only how satellites are built, but, more importantly, how the missions are architected. In fact, this paradigm shift is already underway. As an example, for the SpaceX Falcon9 Transporter missions, one can decide which interface plate to buy and then fit onto that plate a satellite of mass and volume up to the maximum allowed by that plate, at a fixed cost [^19] https://rideshare.spacex.com/search.

### v2

@@ Abstract
Will it be possible in the future to realize large, complex space missions dedicated to basic science like HST, Chandra and JWST? Or will their cost be too great? Today’s space scene is completely different from that of even five years ago, and certainly from that of the time when HST, Chandra and JWST were conceived and built. Space-related investments have grown exponentially in recent years, with a monetary investment exceeding half a trillion dollars per year since 2023. This boom is greatly aided by the rise of the so-called ’new space’ economy driven by private commercial funding, which for the first time last year surpassed public investments in space. The establishment of a market logic to space activities results in more competition and a resulting dramatic cost and schedule reduction. Can space science take advantage of the benefits of the new space economy to reduce cost and development time and at the same time succeed in producing powerful missions in basic science? The prospects for Europe and the United States are considered here. We argue that this goal would be achievable if the scientific community could take advantage of the three pillars underlying the innovation of the new space economy: (1) technology innovation proceeding through both incremental innovation and disruptive innovation, (2) business innovation, through vertical integration, scale production, and service-oriented business model, and (3) cultural innovation, through openness to risk and iterative development.
@@ 1 Basic science from space
    gravitational wave astrophysics and multimessenger astrophysics
@@
The first was still beyond the horizon when the James Webb Space Telescope (JWST, at the time called Next Generation Space Telescope, NGST) was recommended at the end of the 1990s as the highest priority in the 2000 Decadal Survey. The other three were in their infancy, and the first discoveries were just announced. They were not among the main science drivers for NGST. Instead, in the Decadal report ’Astronomy and Astrophysics in the New Millennium’ [1] one reads that NGST ’..is designed to detect light from the first stars and to trace the evolution of galaxies from their formation to the present. It will revolutionize the understanding of how stars and planets form in our galaxy today.. thanks to Kuiper Belt objects (KBOs) in our solar system, the formation of stars and planets in our galaxy, and the dust emission from galaxies out to redshifts of 3’. JWST is doing superb work on the earliest galaxies. Yet JWST is doing a great job on all four of the new themes: up to 40% of JWST time is dedicated to exoplanet studies alone. JWST can be this flexible because it was designed to be a multipurpose observatory that pushes the sensitivity for both photometry and spectroscopy by orders of magnitude with respect to previous or contemporary facilities, opening up new ’discovery space.’ This powerful capability did not come for free. The complexity, cost and development time of JWST are large, even huge, with more than 300 single-point failure modes [2], a cost of about 9B\$ and 23 years elapsed from the publication of the 2000 decadal survey to launch. This cost should be compared with the predicted cost to completion for NGST in ’Astronomy and Astrophysics in the New Millennium’, given as 1B\$. This almost 10-fold cost growth of NGST cannibalized other NASA Astrophysics programs, notably the Beyond Einstein program that spanned a wide range of other wavelengths [3]. Concentration on NGST introduced an enormous risk to space-based astrophysics by putting all its eggs in one basket: A failure of JWST would have meant a decades-long setback for all observational astrophysics.
@@
Even more worrisome is that it is far from certain that NASA, ESA, or any other space agency will be willing or able to embark on a new JWST-scale project (i.e., \$10B, >20>20> 20 year development time). In April 2024 NASA paused the original JPL-led Mars Sample Return project [^2] https://www.nasa.gov/news-release/nasa-sets-path-to-return-mars-samples-seeks-innovative-designs/, which was in the ballpark cost and development time range as JWST. NASA issued a call to industry for alternative approaches and in June 2024 selected seven companies to conduct studies of ’out of the box’ concepts that can deliver samples faster and cheaper than the current plan [^3] https://www.nasa.gov/news-release/nasa-exploring-alternative-mars-sample-return-methods/. This rethought mission may become another example in which NASA buys a service for science-related or exploration-related missions, as NASA now does in other areas, for example, transportation to and from the International Space Station and the Commercial Lunar Payload Service. Another example of the NASA change in attitude is the end of the Moon VIPER rover project [^4] https://www.nasa.gov/news-release/nasa-ends-viper-project-continues-moon-exploration/ due to increases in costs and delays.All this shows that we may have already hit the funding wall [5]. An additional cause for concern to astrophysics is that NASA’s major effort today and for the next few years will certainly be the Artemis project, returning humans to the lunar surface in a long-term program. How can space astrophysics thrive in these circumstances? Can projects like HWO and Artemis coexist? An even greater element of uncertainty is to predict at this time what the attitude of NASA and the new U.S. government in general toward basic science missions might be in the next years. The first signals are definitely not encouraging, given the proposed cuts to the NASA 2026 budget (a two-third cut to NASA astrophysics with respect to 2025, a third cut to planetary science, and a half cut to Earth science [^5] https://www.nasa.gov/wp-content/uploads/2025/05/fy-2026-budget-technical-supplement-002.pdf.
@@
An alternative to expensive multipurpose observatories is to have multiple dedicated missions. ESA’s Gaia mission, for example, opened up new discovery space by measuring the positions and motions of a billion stars some 1000 times better than was possible before. Gaia is becoming the most productive telescope in terms of the number of papers produced from its data, about 2000 papers published per year in refereed journals [^6] https://www.cosmos.esa.int/web/gaia/peer-reviewed-journals, compared with roughly 1000 papers per year based on HST data [^7] https://science.nasa.gov/mission/hubble/overview/hubble-by-the-numbers/. The ESA Euclid and NASA Nancy Grace Roman telescopes are other examples. The total cost of Gaia and Euclid (ESA plus member state contributions) is in the 1-1.5B\$ each, while that of the Roman telescope is close to 5 B\$ (adding a coronagraph to Roman increased its cost by \$ 350-500M [^8] https://roman.gsfc.nasa.gov/science/rsig/2021/Roman_Requirements_20201105.pdf).
Their development time is in the 10-15 years range. That is about 2-10 times cheaper and 2 times faster than JWST and HWO. Although this is an improvement, even at these costs, it is not clear that space agencies can afford more than one per decade of these missions. Such a pace creates the risk that many emerging scientific topics, and so sectors of the astrophysical community, are left unserved at the margin.
@@
Unlike ESA, NASA has a strong Explorer program [^9] https://explorers.gsfc.nasa.gov/ that provides affordable and frequent flight opportunities for smaller science missions [6]. This ensures a rapid response to changing science and technology to enable cutting-edge science at moderate cost. During the long JWST development time, the Explorer program has contributed to maintaining a reasonable level of diversity in the astrophysical ecosystem, avoiding the drying up of entire scientific communities. In Europe, the lack of an Explorer-like program has challenged the high-energy, mid/far-infrared and millimeter communities, which were unable to benefit from replacements for the highly successful XMM-Newton, Herschel and Planck X-ray, far-infrared and CMB-millimeter missions. Of course, Explorer missions are somewhat limited in scope, being dedicated to well-focused scientific goals and in some cases representing case studies and in-orbit demonstrations for new techniques. As an example, the latest SMEX astrophysics mission is IXPE, dedicated to X-ray polarimetry, certainly a niche science. Although IXPE is certainly a highly successful mission (the team was awarded the 2024 Bruno Rossi Prize), its limited collecting area allows sensitive observations of only a few dozen bright sources. IXPE has shown the high potential of X-ray polarimetry, but to fully harvest its fruit, a much larger X-ray polarimetric telescope is needed. The current cost cap of Explorer missions is around \$120M - \$130M for the SMEX missions and about \$300M for MIDEX missions101010\$300M in NASA Fiscal Year (FY) 2022 dollars: https://nspires.nasaprs.com/external/viewrepositorydocument/cmdocumentid=841540/solicitationId=%7B0F29B64B-AFF1-ADAD-7302-75D7AEF2C696%7D/viewSolicitationDocument=1/MIDEX_AO_Amend%201.pdf. From 2000 seven SMEX and five MIDEX astrophysics missions were launched [^11] https://en.wikipedia.org/wiki/Explorers_Program#Small_Explorers_(SMEX), that is about a mission every two years. The typical development time of Explorer missions is 4-5 years, and the entry-level TRL [^12] TRL = technical readiness level: https://en.wikipedia.org/wiki/Technology_readiness_level is 5-6.
@@ 2 The new space economy
At the same time that JWST was launched and HWO was being conceived, we witnessed the booming of the so-called Space Economy. The total amount of money invested in the Space Economy in 2023 was \$570B according to The Space Report of the Space Foundation [^13] https://www.spacefoundation.org/2024/07/18/the-space-report-2024-q2/, with a growth of 7.4% with respect to 2022. According to the Novaspace Space Report 2024 the total size of the Space Economy in 2024 was \$596B  [^14] https://nova.space/hub/product/space-economy-report/, with a further growth of 4.5% compared to the previous year. Recent years have seen private funding overtake public funding. Growth may well accelerate in the next few years, reaching the threshold of \$1000B at the end of this decade and up to 1.8 trillion dollars by 2035151515https://www3.weforum.org/docs/WEF_Space_2024.pdf, dwarfing any conceivable increase in government space agency budgets (except maybe military). Many of these start-ups go well beyond providing launch into on-orbit applications such as Internet communications and many forms of Earth observation. Satellite numbers have grown from less than 200/year in 2000, to almost 3000/year in 2023161616https://www.planet4589.org/space/stats/pay.html .
@@
Can the scientific community take advantage of the new space economy to produce ambitious new space missions dedicated to basic science at a much lower cost and faster timeline? One of the authors, M. Elvis, already discussed what the commercial sector can do for us scientists [5]. Along the same lines, [8] and [9] discuss how space science can benefit from the development of a robust space infrastructure. Here we review all these issues in light of the booming of the new space economy and discuss further questions, such as what the scientific community should do to really exploit the opportunities provided by the commercial space economy. Before addressing these questions, we need to better understand the differences between the old space economy and the new. Table 1 provides a useful simplification as a first approach.
@@
Future: The public-funded space economy will more than double by 2035. Expenditure on defense may triple. Other areas of great increase: PNT; infrastructures. Other areas of public investment for technology development and demonstration: manufacturing in space, including pharmaceutics, biotechnology and semiconductors; space mining; space solar production; space travel; In the past decades activity/business started in the old space-economy framework using public money and then progressed to the new space economy. By 2035 the new space economy may more than triple. Areas of large increases will probably be PNT, which alone may consist of $>$75% of all business, for example last mile deliveries and point-to-point transportation; communications, Earth observation. Other areas can reach maturity and enter the commercial sector. Among them: manufacturing in space, space travel, space mining, space solar production.
@@
Another important space resource will be materials [11, 12]: water, oxygen, hydrogen, construction materials including iron, titanium, and rock, and precious metals, including the Platinum group elements (PGE). Rare Earth elements (REE) are also common on the Moon in the KREEP terrane [13]. REE concentration in the Apollo samples is lower than that of typical REE ores on Earth, but of course the number of samples is limited, and therefore a more detailed prospecting for such elements in promising areas (e.g. Oceanus Procellarum) is desirable. REE were subject to large price volatility in the past decades, in particular when China put strict export quotas in 2009-2010 [14, 15]. Helium 3 concentration on the lunar surface is higher than in most Earth sites, and this element is becoming strategic, for example, to support quantum computers. It is very likely that the first use of materials extracted from the Moon will be in situ, to support astronaut activity (water, oxygen, construction materials) or to produce rocket propellant (hydrogen, oxygen). This will be the first driver of the cis-lunar economy. Only at a later stage it would be conceivable to mine for precious metals or strategic elements that would be brought back to Earth.
The final important resource commonly used in space, but not often acknowledged, is ”knowledge”. Knowledge is crucial to enable or make more efficient harvesting of other resources and to protect space systems. Navigation in low Earth orbits requires detailed knowledge of the Earth’s gravitational field; likewise for the Moon. Navigation near small bodies requires accurate knowledge of their gravitational field and response to solar illumination. Without a deep understanding of the physics of the solar wind and without real-time data, it is impossible to forecast solar storms that can easily damage space infrastructures. And, it is impossible to mine for resources the Moon or Mars or an asteroid if it is not known with precision what is present on the surface or interiors of these bodies through prospecting. These are just a few examples, but it is clear that without long and accurate observations, big data analysis, and the development of accurate models it is impossible to imagine any activity in space. Knowledge therefore has its own very concrete value as intellectual property (IP) in the new space economy. This is a major change from the open data policies of most government space agencies.
@@ 2.1 Innovation paradigms
    Innovation in technology, proceeding through both incremental innovation and disruptive innovation. Disruptive innovation starts with low performance and high cost, but quickly overtakes mainstream technologies. The disruptors then engage in incremental innovation, beginning a new cycle.
@@
    Cultural innovation. Startups must have a bold risk culture, which manifests itself in risk openness project management models, for example the Agile methodology, which breaks projects into several ’sprints’ and delivers a Minimum Viable Product (MVP) at the end of each sprint.
@@ 2.2 Costs: launch
Figure 1, adapted from Visual Capitalist [^18] https://www.visualcapitalist.com/the-cost-of-space-flight/, shows the cost per kg to orbit as a function of time. Due to reusability, the SpaceX Falcon 9 (F9) and Falcon Heavy (FH) launchers have lowered the cost to orbit offered to third parties by a factor of 3-6 with respect to other Western, Russian and Chinese launchers. F9 capacity to low Earth orbit (LEO) is about 20 tons, at an advertised cost of about 62M\$, that is about \$3000/kg.
@@
Today, F9 has an effective monopoly on the launch market. Forthcoming competition with Blue Origin New Glenn may result in a lowered cost for third parties. Table 3 provides examples of launch costs for recent scientific missions. None approach the \$1-3k/kg that is achievable with F9, mainly because they were unable to make use of the full F9 capacity on dedicated launches, especially for missions not designed for an F9 or FH launch. Other causes of high-billed costs include requests for orbits more demanding than the LEO and absence of competition.
@@
Mission Mass kg Orbit Launcher Cost M\$ Cost k\$/kg
@@
The launch cost of Figure 1 has been reached for SpaceX Starlink satellites that have been estimated at 0.7-1.4 M\$ per Starlink 2 mini satellite, with a cost per kg as low as \$1000-\$2000/kg. This should be regarded as the cost of launch for optimized payloads, net of profit. SpaceX claims that the coming on-line of Starship in the next few years will lower the cost by another factor of 10 with respect to F9. The Starship revolution will not be limited to cost to orbit. Starship performance will make it unnecessary to spend time and resources to minimize mass and size [7, 16, 17, 18], which will change not only the way satellites are built, but, more importantly, the way missions are architected. In fact, this paradigm shift is already underway. As an example, for SpaceX Falcon9 Transporter missions, one can decide which interface plate to buy and then fit on that plate a satellite of mass and volume up to the maximum allowed by that plate, at a fixed cost [^22] https://rideshare.spacex.com/search.